OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

jhonyy9 (jhonyy9):

opinion ?

OpenStudy (anonymous):

x=-1

jhonyy9 (jhonyy9):

3x+3=-2x+2-4 can you continue ?

OpenStudy (anonymous):

you had most of the answer try doing it

OpenStudy (anonymous):

thank you @surajpatil083! :)

OpenStudy (anonymous):

3x+3=-2x+2-4 3x+2x=2-4-3 5x=-5 x=-1